Two microphones are placed 68 cm apart on the same side of a loudspeaker And colimear with it the microphones are connected to the two channels of an ascilloscope for a loudspeaker sound of frequency 3000hz the 2 vibrations observed on the screeen of oscilloscope are in phase the frequency is decreased and the vibration are in phase again for frequency 2500hz calculate the speed of sound in air
1
Expert's answer
2018-10-03T06:48:09-0400
l=nv/f_1 =(n-1)v/f_2
0.68 =nv/3000 =(n-1)v/2500
Thus,
2500n=3000(n-1)
500n=3000
n=3000/500=6.
The speed of sound in air:
v=0.68(3000)/6=340 m/s.